Grade 11Algebra

Find equation of circle which touches X+2y=3 at (1,1) and it also touches x-axis?

Rohan patil
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
Let the centre = (a, b)
 
Hence
 
Sqrt [(a-1)² + (b -1)²] =( a +2b -3 )/sqrt5 = b = radius
 
Now you can solve this by taking twi equations together